OpenStudy (anonymous):

To evaluate this limit: lim sqroot(4-x)-2 x-0 ------ x Could you not simply re-arrange the numerator to get rid of everything except the x, since sqroot4-2 would equal 0, leaving the -x over x?

OpenStudy (anonymous):

The example in the book says I must conjugate, but this seems simpler if there is a way to get around it ending up in -x/x.

ganeshie8 (ganeshie8):

\[\large \lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x}\] like this ?

OpenStudy (anonymous):

Yes, exactly.

ganeshie8 (ganeshie8):

I'm still wondering how do you think the numerator can be simlified to \(\large -x\) without doing the conjugate thingy hmm

ganeshie8 (ganeshie8):

Notice entire \(\large 4-x\) is inside squareroot not just \(\large 4\)

OpenStudy (anonymous):

Must they be taken as one number then? It cannot be both √4 and √-x?

ganeshie8 (ganeshie8):

you cannot break squareroot like that

OpenStudy (anonymous):

Ah, my mistake then

ganeshie8 (ganeshie8):

\[\sqrt{a-b} \ne \sqrt{a}-\sqrt{b}\]

OpenStudy (anonymous):

I assume I must conjugate to evaluate this then?

ganeshie8 (ganeshie8):

thats right but how does a conjugate help here and why is it required here ?

ganeshie8 (ganeshie8):

why can't we simply plugin x = 0 ?

OpenStudy (anonymous):

Would it not then be indeterminate?

ganeshie8 (ganeshie8):

Exactly! so we try to change the form by multiplying it by a special kind of 1 : \[1 = \dfrac{\sqrt{4-x}+2}{\sqrt{4-x}+2}\]

ganeshie8 (ganeshie8):

that magically cancels the bad stuff and allows you to plugin x = 0 and get a determinate form

OpenStudy (anonymous):

The book I am using teaches this method as conjugation, though I've never seen this term used in this manner before.

ganeshie8 (ganeshie8):

\[\large \begin{align} \lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x} &= \lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x}\times \dfrac{\sqrt{4-x}+2}{\sqrt{4-x}+2} \\~\\ &= \cdots \end{align}\]

ganeshie8 (ganeshie8):

yes "conjugate" is a name which has similar/different meanings in different contexts here conjugate just refers to changing the sign of middle term in a binomial

ganeshie8 (ganeshie8):

for example conjugate of \(2-\sqrt{3}\) is \(2+\sqrt{3}\)

OpenStudy (anonymous):

I am currently stuck on reconciling the denominator \[x(\sqrt{4-x}+2)\] to be the same as the numerator \[\sqrt{4-x}^2-2^2\]

OpenStudy (anonymous):

At least, I assume that is the next step.

ganeshie8 (ganeshie8):

you're doing good with the numerator, keep going..

OpenStudy (kainui):

The good old "Let's multiply by a fancy form of 1" trick to change your perspective to make the problem to be easier strikes yet again. =D

OpenStudy (anonymous):

Now I have\[\frac{ 4-x-4 }{ }\], but still no idea how to get the denominator to match

OpenStudy (anonymous):

And I'm not too sure of the end result I've come up with on the numerator as well

ganeshie8 (ganeshie8):

:) have faith and keep going..

OpenStudy (anonymous):

I guess it would be 4-x+4. since -2^2 is +4...

OpenStudy (anonymous):

no,it's not (-2)^2 it's -2^2

ganeshie8 (ganeshie8):

4-x-4 simplifies to -x yes ? next look at denominator, there is a x waiting to get cancelled with the numerator x

OpenStudy (anonymous):

But the \[\sqrt{4-x}\] cannot be simplified with just the x, can it?

OpenStudy (anonymous):

\[\lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x} = \lim_{x \rightarrow 0} \frac{ -1 }{ \sqrt{4-x} + 2 }\]

OpenStudy (anonymous):

now what's ur idea? do u think you can just put x=0 in that and find it?

OpenStudy (anonymous):

I apologize but I cannot quite see where I'm going wrong here.

OpenStudy (anonymous):

@ganeshie8 , sry u were helping out ;(

OpenStudy (anonymous):

@Jaeriko , what have u got so far?

OpenStudy (anonymous):

At the moment, my equation looks like this: \[\frac{ \sqrt{4-x}^2 -2^2}{ x(\sqrt{4-x } +2)}\]

OpenStudy (anonymous):

correct ;)

OpenStudy (anonymous):

Unsure of whether the numerator will simplify to -x or x, and completely in the dark as to how to simplify the denom.

OpenStudy (anonymous):

I do not understand the difference between -2^2 and (-2)^2, for the numerator issue.

OpenStudy (anonymous):

As both are -2 by -2, as I understand it.

OpenStudy (anonymous):

well, \(-a ^{2} = (-1)(a^{2})\)

ganeshie8 (ganeshie8):

thats a good question

OpenStudy (anonymous):

u see in that equation we have \( -2^2 \)

OpenStudy (anonymous):

it means you should first find \( 2^2 \) then make it negative.

OpenStudy (anonymous):

I understand what you're saying there, I've just never heard of it used like that before.

OpenStudy (anonymous):

Nor have I ever run into an equation where it was not simply (-x)x(-x) in a -x^2 situation. Is there a name for this rule I can look up?

OpenStudy (anonymous):

no i havn't heard name of a rule for it...you know i DIDN'T catch what ur saying.please state WHAT you have problem with...you don't know how you can find the numerator?

ganeshie8 (ganeshie8):

let me ask you a question @Jaeriko whats the value of \(\large 3^2 - 2^2\) ?

OpenStudy (anonymous):

I do not understand how I can simplify the current state of my equations numerator to the -1 that you and the book state is the solution.

OpenStudy (anonymous):

5, if I understand correctly.

ganeshie8 (ganeshie8):

Right! how did you get 5 ?

OpenStudy (anonymous):

Ah, I see

OpenStudy (anonymous):

I suppose I've done this many times without truly noticing, then

ganeshie8 (ganeshie8):

\[3^2 - 2^2 = 9 - 4 = 5\] similarly \[(\sqrt{4-x})^2 - 2^2 = 4-x - 4 = -x\]

OpenStudy (anonymous):

@ganeshie8 nice explanation ;)

ganeshie8 (ganeshie8):

its bit hard to explain.. but i see you figured it out already :)

OpenStudy (anonymous):

Okay, I understand that step now...but what about turning -x into -1?

OpenStudy (anonymous):

Or does that come later?

ganeshie8 (ganeshie8):

that comes later -x wont change into 1 just like that without some balck magic haha!

OpenStudy (anonymous):

well \(\frac{ -x }{ x } = -1\)

OpenStudy (anonymous):

Okay, so how do we get eliminate the \[x( \sqrt{4-x}+2)\] on the bottom now? I don't know how we can do anything with the \[\sqrt{4-x}\], as we have to deal with them as a group instead of individual numbers.

OpenStudy (anonymous):

Can we have \[\sqrt{4-x^2}\] now?

ganeshie8 (ganeshie8):

its much simpler than that the only bad guy in the denominator is \(x\) because it pulls the denominator to 0

ganeshie8 (ganeshie8):

just cancel the x in denominator with the numerator x then plugin x = 0 you're done!

ganeshie8 (ganeshie8):

here everything at one place : \[\require{cancel}\large \begin{align} \lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x} &= \lim\limits_{x\to 0} \, \dfrac{\sqrt{4-x}-2}{x}\times \dfrac{\sqrt{4-x}+2}{\sqrt{4-x}+2} \\~\\ &= \lim\limits_{x\to 0} \,\dfrac{(\sqrt{4-x})^2 - 2^2}{x(\sqrt{4-x}+2)} \\~\\ &= \lim\limits_{x\to 0} \,\dfrac{4-x-4}{x(\sqrt{4-x}+2)} \\~\\ &= \lim\limits_{x\to 0} \,\dfrac{-x}{x(\sqrt{4-x}+2)} \\~\\ &= \lim\limits_{x\to 0} \,\dfrac{-\cancel{x}}{\cancel{x}(\sqrt{4-x}+2)} \\~\\ &= \lim\limits_{x\to 0} \,\dfrac{-1}{\sqrt{4-x}+2} \\~\\ &=\dfrac{-1}{\sqrt{4-0}+2} \\~\\ &=\dfrac{-1}{2+2} \\~\\ &=\dfrac{-1}{4} \end{align}\]

OpenStudy (anonymous):

Oh! That's amazing, I didn't even see it coming for some reason. Thank both of you so much for your help!

OpenStudy (anonymous):

As soon as I realized you could cancel the -x and x with each other to leave the -1, it feel into place.

OpenStudy (anonymous):

I wish I had but more medals to give

ganeshie8 (ganeshie8):

yes you get use to these after working one two examples and will feel its really easy yw :)

OpenStudy (anonymous):

ur welcome @ganeshie8 helped you well for his awesome explanation.... @Jaeriko wait a moment,i have prepared some awesome problems,you can do them for fun and check your understanding

OpenStudy (anonymous):

\(\large \lim_{h \rightarrow 0} \frac{ \sqrt{1+h} - 1}{ h }\) \(\large \lim_{x \rightarrow -4} \frac{ \sqrt{x^2 + 9} -5 }{ x+4 }\) \(\large \lim_{x \rightarrow 16} \frac{ 4-\sqrt{x} }{ 16x - x^2}\) \(\large \lim_{x \rightarrow 7} \frac{ \sqrt{x+2}-3 }{ x-7 }\)

OpenStudy (anonymous):

I greatly appreciate the time you have both spent explaining this to me. Have a good night, and thank you again. @PFEH.1999 ,I'm writing those down for practice after I finish work tomorrow. Thank you so much for your time and effort!

ganeshie8 (ganeshie8):

Nice :) All of these can be worked out using conjugate thingy.. also they all can be evaluated by using definition of derivative !